OpenStudy (aroub):

Evaluate lim x->pi/6 (sqrt3 cosx)

OpenStudy (aroub):

\[(\sqrt{3} \cos x)\]

OpenStudy (anonymous):

Does cos(x) also shares room with 3 in that square root??

OpenStudy (anonymous):

Okay.. I got it..

OpenStudy (aroub):

Oh, wait! You only substitute?

OpenStudy (aroub):

No, it doesnt

OpenStudy (anonymous):

I am looking for what is a bit tricky part here..

OpenStudy (aroub):

lol

OpenStudy (aroub):

There isnt any eh?

OpenStudy (anonymous):

Just put x = pi/6 there.. and you are done..

OpenStudy (anonymous):

Till you are getting a finite value on substituting, you can substitute the values directly..

OpenStudy (aroub):

Yeah, i just got that xD

OpenStudy (anonymous):

If cos(x) is under root, then also, you will substitute directly..

OpenStudy (aroub):

Yes, exactly :)

OpenStudy (anonymous):

Good.. :)

OpenStudy (anonymous):

Should I give you one??

OpenStudy (aroub):

That would be great!

OpenStudy (anonymous):

Okay: if you want to solve it here, then you can solve it otherwise you can make it your question too and post it also, I leave it on you.. :)

OpenStudy (anonymous):

Try: \[\lim_{x \to -2} \frac{(x^2 - 4)}{(x+ 2)}\]

OpenStudy (anonymous):

x is tending to -2 there, be careful there..

OpenStudy (aroub):

-4 :D

OpenStudy (anonymous):

Oh..!! I am not good at framing questions... :P

hartnn (hartnn):

let me give one

OpenStudy (anonymous):

@hartnn you wanna try??

hartnn (hartnn):

limit x->0 of (sin x+x)/ tan x

OpenStudy (aroub):

No, you're good, but I'm smart :p

OpenStudy (anonymous):

Hardik, keep it somewhat simpler one, as I have seen she or he is doing basics of limits.. :)

hartnn (hartnn):

it is simple fr smart people ;)

hartnn (hartnn):

and she also know the sin x/x formula!

OpenStudy (anonymous):

And she also knows about tan(x)/x, I just told her.. :P

hartnn (hartnn):

:P

OpenStudy (anonymous):

@aroub just take your time, if you have doubt anywhere and then let us know, you don't worry about anything if you are not getting. But I think, just a small use of your brain will take you there..

OpenStudy (aroub):

Now that's a tricky one!

OpenStudy (aroub):

is it 1?

hartnn (hartnn):

nopes, try again,

hartnn (hartnn):

how you got one ? will help u spot the error

OpenStudy (aroub):

Haha, okay!

hartnn (hartnn):

whenever you see sin x you should think, how can i use the formula lim x->0 sin x/x =1

hartnn (hartnn):

or tan x because tan x is sin x/cos x and cos 0 =1

OpenStudy (aroub):

Thats what i'm doing

hartnn (hartnn):

how would u get sin x/x and tan x/x ?

OpenStudy (aroub):

just the "+x" is annoying

hartnn (hartnn):

you have sin x in the numerator, how would u get the sin x/x ?

OpenStudy (aroub):

okay, wait, let me do this again and i'll come back to you :)

hartnn (hartnn):

sure :)

OpenStudy (aroub):

Do you get to something like this: \[\lim x \rightarrow0 \frac{ sinx+x }{ sinx }\]

hartnn (hartnn):

just replaced tan x by sin x ? the limiting value will be same as cos 0=1

OpenStudy (aroub):

No, its actually like this: \[\lim_{x \rightarrow 0} \frac{ sinx+x }{ sinx } \lim_{x \rightarrow 0} cosx\]

OpenStudy (aroub):

but yeah, you're right

hartnn (hartnn):

yes, your step is correct :) so now try to find lim x->0 (sin x+x)/sin x

OpenStudy (aroub):

Can I separate them? like this: sinx/sinx + x/sinx

hartnn (hartnn):

sure you can! (there is an easier way, but you can do that too)

hartnn (hartnn):

now how will you find lim x->0 x/sin x ?

OpenStudy (aroub):

If i flip (do the reciprocal) everything does the limit change?

OpenStudy (aroub):

please tell me its 2? :p

hartnn (hartnn):

ofcourse it is 2 ! sorry for the suspense :P

hartnn (hartnn):

if lim f(x) = a then lim 1/f(x) = 1/ lim f(x) = 1/a

OpenStudy (aroub):

Hahah, finally! Thank youuuuuu :D

hartnn (hartnn):

most welcome ^_^ now the easier way

hartnn (hartnn):

(sin x +x)/ tan x = (sin x/x + x/x)/ (tan x/x) this i got by dividing 'x' in the numerator and denominator that i did because i had sin x and wanted sin x/x

hartnn (hartnn):

after that step, its a piece of cake sin x/x = 1, x/x = 1 tanx /x =1

OpenStudy (aroub):

That's very funny actually compared to what I did xD

hartnn (hartnn):

but i am glad you tried it on your own, now you know 2 methods to solve that :)

OpenStudy (anonymous):

It can become very simple if you have separated it when you first saw it.. :) \[\lim_{x \to 0} (\frac{\sin(x)}{\tan(x)} + \frac{x}{\tan(x)}) \implies \lim_{x \to 0}(\cos(x) + \frac{x}{\tan(x)}) \\ \lim_{x \to 0} (\cos(x)) + \lim_{x \to 0}(\frac{x}{\tan(x)})\]

OpenStudy (anonymous):

Now you must know that: \[\lim_{x \to 0} (\frac{\tan(x)}{x}) = \lim_{x \to 0}(\frac{x}{\tan(x)}) = \color{blue}{1}\]

OpenStudy (aroub):

I would've never thought of that haha :p You can also flip it from the beginning lim x->0 tanx/sinx+x or multiply sinx to both the numerator and the denominator so there's kinda 3 methods on solving this :p

OpenStudy (anonymous):

Also : \(cos(0) = 1\)..

hartnn (hartnn):

there are many methods to solve every math problem :3

OpenStudy (aroub):

4 lol

OpenStudy (aroub):

Yes, that's true!

OpenStudy (anonymous):

Actually, it is 1 method only, the thing is you are going to once final while playing in between.. :P

OpenStudy (anonymous):

*one.

hartnn (hartnn):

move onto next question :P don't play with this on, the entire day :P

hartnn (hartnn):

*one

OpenStudy (anonymous):

*one final place.. :P

OpenStudy (aroub):

Haha, yes! I'm off to bed anyway. Thank you both so much!

OpenStudy (anonymous):

yeah, I am unable to see myself in open list of questions on my left, because we are wandering in closed section of openstudy..

OpenStudy (anonymous):

Good Night.. :)